The heat liberatedwhen 1.89g of benzoic acid is burnt in a bomb calorimeter at 298K increases the temp of 18.94Kg of water by 0.632K.If specific heat of water at 298K is 0.998 cal/gK,the value of heat of combustion of benzoic acid is what???????? Ans-852.7kcal HOW??????????

Grade:12th pass

1 Answers

sakshi aggarwal
5 years ago
Given, Weight of benzoic acid = 1.89 gTemperature of bomb calorimeter = 25∘C25∘CMass of water (m) = 18.94 kg = 18940 gIncrese in temperature (ΔTΔT) = 0.632∘C0.632∘Cans Specific heat of water (s) = 0.998 cal/g/degWe know that heat gained by water or heat liberated by benzoic acid (Q) = msΔT=18940×0.998×0.632=11946.14calmsΔT=18940×0.998×0.632=11946.14calSince, 1.89 g of acid liberates 11946.14 cal of heat,∴∴ heat liberated by 122 g of acid (1 mol) = 11946.14×1221.8911946.14×1221.89 = 77126.5 cal = 771.12 kcal(where 122 g is the molecular weight of benzoic acid)